OpenStudy (anonymous):

what is the equation of the line that has a slope of 4 and passes through the point (3, -10)

OpenStudy (anonymous):

Sierra, you need to use the 'point-slope' version of a straight line to work out your equation. That is,\[y-y_1=m(x-x_1)\]where m is your slope and (x_1,y_1) is your given point. So, for you,\[y-(-10)=4(x-3) \rightarrow y+10=4x-12\]so\[y=4x-2\]

OpenStudy (anonymous):

u mean y = 4x - 22 right? and thank u sooo much!

OpenStudy (anonymous):

Woops, yes, I do. Didn't tap the '2' twice!

OpenStudy (anonymous):

haha thank u!

OpenStudy (anonymous):

np